# Model Equality Testing: Which Model Is This API Serving?

### v1

@@ Results.
We now ask how well the Hamming MMD correlates with task accuracy when available. Automated evaluation exists for one of our prompt distributions, HumanEval. We find a moderate positive correlation between the absolute average accuracy difference and the Hamming MMD (Figure 4 right, $R^{2}=0.392$). In several cases, the MMD is high but the accuracy difference is low. These are often when both task accuracies are low: the gray points in the figure highlight pairs where both distributions have accuracy $<$ 10%. In these cases, the MMD captures that there are multiple ways to be wrong for a task. For example, the bottom-right-most point in the figure contrasts Llama-3.1 8B from Perplexity vs. the nf4 version; Perplexity answers with incorrect code snippets, while nf4 outputs nonsense completions (e.g., “simulation!!!!!!!!!![…]”). Samples from this comparison can be found in Appendix D.3.
@@ 6 Related work
Auditing ML APIs. Prior work audits APIs by monitoring benchmark performance. Most works focus on classifier APIs (Chen et al., 2021a; 2022), while work on language models has focused only on low-temperature decoding with multiple choice or short-answer tasks (ArtificialAnalysis, 2024; Chen et al., 2023; Eyuboglu et al., 2024). In contrast, we take a hypothesis testing approach to auditing language model APIs, emphasizing the sampling setting, and we quantify distances between string samples. This approach allows us to audit APIs on tasks that do not have crisp automatic evaluation metrics. In cryptography, a body of work has aimed to verify ML API predictions through proof protocols (Ghodsi et al., 2017; Feng et al., 2021; Liu et al., 2021; Kang et al., 2022; Weng et al., 2023; Lee et al., 2024). These methods require APIs to opt-in to providing proofs of valid inference alongside each prediction; public verifiers then check these proofs for correctness with perfect accuracy. Unfortunately, these methods require ongoing cooperation from the API provider, and they scale too poorly to apply to large language models: in Kang et al. (2022), generating a proof takes 41 minutes per prediction for a 68M parameter MobileNet (Sandler et al., 2018).
@@ Acknowledgements
PL is a co-founder of Together AI; this work was done in his Stanford capacity. The topic, research, and results of this work were not shared with Together, or any other API provider evaluated, until the public release of the paper. IG and CG have no conflicts of interest with respect to the providers or developers evaluated in this paper.
@@ A.2 Simulating p-values
    Simulating the test statistic under the null (Algorithm 1). This is done by repeatedly sampling $\mathcal{D}_{Q}$ and $\mathcal{D}_{P}$ from $P$ and caching $\widehat{\text{MMD}}(\mathcal{D}_{Q},\mathcal{D}_{P})$. The p-value is then the proportion of times the test statistic is greater than or equal to the observed test statistic. We conduct tests using this method in the main text, reusing the same cached empirical distribution of the test statistic under the null for all alternatives at that sample size. Note that this method requires significant sampling access to $P$.
@@
7:Compute p-value as the proportion of $\mathcal{T}$ greater than or equal to $\widehat{\text{MMD}}(\mathcal{D}_{Q},\mathcal{D}_{P})$
@@
8:Compute p-value as the proportion of $\mathcal{T}$ greater than or equal to $\widehat{\text{MMD}}(\mathcal{D}_{Q},\mathcal{D}_{P})$
@@ Appendix C Additional results
C.1 Sample-complexity and length
This appendix includes additional results from §4.1. Figures 5, 6, and 7 stratify the sample complexity and completion length results by the alternative distribution $Q$ and model.
Table 8 shows the power of the Hamming MMD test to distinguish between pairs of models (and other alternatives) in $L=$1000$$ character space. In general, moving to this higher-dimensional space decreases the power of the test.
Figure 5: Sample complexities for different two-sample tests. Each subplot represents power for an alternative distribution $Q$ averaged over five language models and ten prompt distributions $\pi$. The Hamming MMD test is significantly more sample-efficient for all alternatives except fp16 quantization: for this alternative, while other two-sample tests can attain perfect power with enough samples, the kernel test increases power slowly. This suggests that fp16 and fp32 differ in ways that are difficult to capture with the Hamming kernel.
@@ C.1 Sample-complexity and length
Figure 6: Sample complexities for different two-sample tests. Each subplot represents power for a particular alternative distribution $Q$ and model, but averaged over ten prompt distributions $\pi$. Some model and alternative combinations are more difficult to detect than others.
@@
Figure 7: Simulated powers for different completion lengths $L$, stratified by the alternative distribution $Q$. Across all alternatives, kernel tests suffer less drop in power as the length of completions increases. Traditional two-sample tests observe a slight power increase from $L=1$ tokens to $L=5$, tokens, but power dramatically degrades afterwards.
Mistral 7B 3 8B 3.1 8B 3 70B 3.1 70B
Mistral 7B 0.06 1.00 1.00 1.00 1.00
Llama-3 8B 1.00 0.05 0.76 0.98 0.95
Llama-3.1 8B 1.00 0.83 0.07 1.00 0.53
Llama-3 70B 1.00 0.99 1.00 0.07 0.89
Llama-3.1 70B 1.00 0.98 0.75 0.99 0.06
Phi-3 mini 1.00 1.00 1.00 1.00 1.00
Gemma-2 9B 1.00 1.00 1.00 1.00 1.00
OLMo 7B 1.00 1.00 1.00 1.00 1.00
GPT-4o mini 1.00 1.00 1.00 1.00 1.00
gpt-4o 1.00 1.00 1.00 1.00 1.00
gpt-3.5-turbo-0125 1.00 1.00 1.00 1.00 1.00
gpt-3.5-turbo-1106 1.00 1.00 1.00 1.00 1.00
gpt-4-0125-preview 1.00 1.00 1.00 1.00 1.00
watermark 0.23 0.62 0.32 0.57 0.26
int8 0.15 0.30 0.07 1.00 0.99
nf4 0.44 0.38 1.00 1.00 1.00
@@
Table 8: Power of the Hamming MMD test to distinguish between pairs of models (and other alternatives) in $L=$1000$$ character space. The column represents the null distribution $P$. Italicized entries are FPRs. Note that unlike §4.1, the test is now run in higher-dimensional character space, which generally reduces power.
C.2 Effect of the prompt distribution
@@
In §4.1, we evaluate tests on a Wikipedia prompt distribution with support over $m=25$ prompts. Here, we evaluate the effect of $m$ on the MMD Hamming test’s power, with $N=10m$ fixed to an average of 10 samples per prompt, and $L=50$ tokens (Figure 8). Larger $m$ values increase the power of the test, suggesting that users benefit from testing many prompts together, so long as the sample size is increased proportionally.
@@
C.3 Extension: comparing two-sample and approximate goodness-of-fit tests
@@
C.4 Extension: asymmetric sampling costs

### v2

@@ Results.
We now ask how well the Hamming MMD correlates with task accuracy when available. Automated verifiers exist for one of our prompt distributions, HumanEval. We find a moderate positive correlation between the absolute average accuracy difference and the Hamming MMD (Figure 4 right, $R^{2}=0.392$). In several cases, the MMD is high but the accuracy difference is low. These are often when both task accuracies are low: the gray points in the figure highlight pairs where both distributions have accuracy $<$ 10%. In these cases, the MMD captures that there are multiple ways to be wrong for a task. For example, the bottom-right-most point in the figure contrasts Llama-3.1 8B from Perplexity vs. the nf4 version; Perplexity answers with incorrect code snippets, while nf4 outputs nonsense completions (e.g., “simulation!!!!!!!!!![…]”). Samples from this comparison can be found in Appendix D.3.
@@ 6 Related work
Auditing ML APIs. Prior work either only applies to the classifier setting (Chen et al., 2021a; 2022) or audit by monitoring performance on low-temperature multiple-choice or short-answer benchmarks (ArtificialAnalysis, 2024; Chen et al., 2023; Eyuboglu et al., 2024). Concurrent work studies a general language model setting, but does not provide guarantees of audit correctness (Sam et al., 2025). In contrast, we formalize auditing as a hypothesis testing problem, allowing us to provide FPR guarantees, and we study the general setting of sampling completions from language models, including for tasks without automated verifiers. In cryptography, a body of work has aimed to verify ML API predictions through proof protocols (Ghodsi et al., 2017; Feng et al., 2021; Liu et al., 2021; Kang et al., 2022; Weng et al., 2023; Lee et al., 2024). These methods require APIs to cooperate and opt-in to providing proofs of valid inference alongside each prediction; public verifiers then check these proofs for correctness with perfect accuracy. Unfortunately, these methods scale too poorly to apply to large language models: in Kang et al. (2022), generating a proof takes 41 minutes per prediction for a 68M parameter MobileNet (Sandler et al., 2018).
@@ Acknowledgements
PL is a co-founder of Together AI; this work was done in his Stanford capacity. The topic, research, and results of this work were not shared with Together, or any other API provider evaluated, until the public release of the paper.
@@ A.2 Simulating p-values
    Simulating the test statistic under the null (Algorithm 1). This is done by repeatedly sampling $\mathcal{D}_{Q}$ and $\mathcal{D}_{P}$ from $P$ and caching $\widehat{\text{MMD}}(\mathcal{D}_{P},\mathcal{D}_{Q})$. The p-value is then the proportion of times the test statistic is greater than or equal to the observed test statistic. We conduct tests using this method in the main text, reusing the same cached empirical distribution of the test statistic under the null for all alternatives at that sample size. Note that this method requires significant sampling access to $P$.
@@
7:Compute p-value as the proportion of $\mathcal{T}$ greater than or equal to $\widehat{\text{MMD}}(\mathcal{D}_{P},\mathcal{D}_{Q})$
@@
8:Compute p-value as the proportion of $\mathcal{T}$ greater than or equal to $\widehat{\text{MMD}}(\mathcal{D}_{P},\mathcal{D}_{Q})$
@@ Appendix C Additional results
C.1 Additional results from §4.1
In §4.1, we evaluated two-sample tests on their ability to distinguish samples (in token space, $L\leq 50$) from full-precision models from those of quantized models and watermarked models. Concretely, we evaluated tests’ power on pairs of distributions $(P,Q)$, where $P$ represents samples from the fp32 model and $Q$ represents samples from one of {fp16, int8, nf4, or watermarked} versions of the same model. To ensure generalizable results, we experimented with five language models (Mistral 7B Instruct, Llama-3 8B and 70B Instruct, and Llama-3.1 8B and 70B Instruct) and ten Wikipedia-based prompt distributions (Box 4).
Figure 2 in the main text plotted sample complexity and length results, averaging across language models and prompt distributions. In this appendix, we show plots for the same experiments, but stratified by language model and prompt distribution.
  • Figure 5 shows the sample complexities for different two-sample tests, stratified by the alternative distribution $Q$, but averaged over the five language models and ten prompt distributions.
@@
  • Figure 6 shows the sample complexities for different two-sample tests, stratified by the alternative distribution $Q$ and model, but averaged over ten prompt distributions.
  • Figure 7 shows the simulated powers for different completion lengths $L$, stratified by the alternative distribution $Q$.
Additionally, Table 8 provides the power of the Hamming MMD test to distinguish $(P,Q)$ for other language models in $L=50$ token space. This table adds Phi-3 Mini (Instruct), OLMo 7B (Instruct), and Gemma-2 9B (Instruct). Table 9 evaluates a Levenshtein MMD test, which uses the Levenshtein distance as the kernel function. The Levenshtein kernel is defined as $k(x,y)=\exp(-\text{Levenshtein}(x,y))$. It is less powerful than the Hamming kernel, but still outperforms traditional two-sample tests.
Discussion: the Hamming MMD is the most powerful test.
Across Figure 2 (left), Figure 5, and Figure 6, we observe that the Hamming MMD test is significantly more sample-efficient for all alternatives except fp16 quantization at completion length $L=50$ tokens. We also observe that the Hamming MMD is more robust to changes in completion length than other tests (Figure 2 middle, Figure 7).
Discussion: tests’ powers depend on the particular $(P,Q)$ pair.
In Figure 6 and Table 8, we observe that tests’ powers are heterogeneous across different $(P,Q)$ pairs. The choice of model affects both the reference distribution $P$ and the effect of interventions like quantization, i.e., the alternative $Q$. For example, 4-bit quantization of the Llama and OLMo models is consistently noticeable. On the other hand, 8-bit quantization and watermarking have more inconsistent effects. The impact of quantization is heterogeneous even within the same model family: for example, the Llama-3 8B model is more affected by int8 quantization than the Llama-3.1 8B model. On average however, as discussed above, our results above still support our claim that the Hamming MMD is more powerful than other kernel choices.
Figure 5: Sample complexities for different two-sample tests. Each subplot represents power for an alternative distribution $Q$ averaged over five language models and ten prompt distributions $\pi$. Tests are conducted with sample sizes ranging from $N=10m$ to $N=100m$, where $m=25$ is the number of prompts in the prompt distribution. The completion length is fixed to $L=50$ tokens. The Hamming MMD test is significantly more sample-efficient for all alternatives except fp16 quantization: for this alternative, while other two-sample tests can attain perfect power with enough samples, the kernel test increases power slowly. This suggests that fp16 and fp32 differ in ways that are difficult to capture with the Hamming kernel.
Figure 6: Sample complexities for different two-sample tests. Each subplot represents power for a particular alternative distribution $Q$ and model, but averaged over ten prompt distributions $\pi$. Tests are conducted with sample sizes ranging from $N=10m$ to $N=100m$, where $m=25$ is the number of prompts in the prompt distribution. The completion length is fixed to $L=50$ tokens. Some model and alternative combinations are more difficult to detect than others.
@@
Figure 7: Simulated powers for different completion lengths $L\in\{1,\cdots,50\}$, stratified by the alternative distribution $Q$. The sample size is fixed to $N=10m$, where $m=25$ is the number of prompts in the prompt distribution. Across all alternatives, kernel tests suffer less drop in power as the length of completions increases. Traditional two-sample tests observe a slight power increase from $L=1$ tokens to $L=5$, tokens, but power dramatically degrades afterwards.
Model watermark nf4 int8 FPR (fp32)
Llama-3 8B 1.00 (0.00) 0.93 (0.04) 0.79 (0.08) 0.05 (0.01)
Llama-3.1 8B 0.62 (0.04) 1.00 (0.00) 0.08 (0.02) 0.05 (0.01)
Llama-3 70B 0.97 (0.01) 1.00 (0.00) 1.00 (0.00) 0.07 (0.01)
Llama-3.1 70B 0.54 (0.06) 1.00 (0.00) 1.00 (0.00) 0.06 (0.01)
Mistral 7B 0.72 (0.06) 0.89 (0.04) 0.30 (0.06) 0.04 (0.01)
OLMo 7B 0.47 (0.08) 0.99 (0.01) 0.36 (0.08) 0.06 (0.01)
Gemma-2 9B 0.43 (0.02) 0.12 (0.01) 0.06 (0.01) 0.06 (0.01)
Phi-3 Mini 0.75 (0.04) 0.63 (0.05) 0.24 (0.01) 0.06 (0.01)
Table 8: Power of the Hamming MMD test to distinguish the fp32 model from alternatives in $L=50$ token space at $N=10m$ samples, where $m=25$ is the number of prompts in the prompt distribution. Averages and standard errors are reported over ten prompt distributions. The column represents the alternative distribution $Q$. The FPR (fp32) column represents the false positive rate when comparing the full-precision model to itself.
Model watermark nf4 int8 FPR (fp32)
Llama-3 8B 0.14 (0.01) 0.15 (0.01) 0.16 (0.01) 0.07 (0.01)
Llama-3 70B 0.13 (0.02) 1.00 (0.00) 0.12 (0.01) 0.04 (0.01)
Mistral 7B 0.11 (0.01) 0.19 (0.01) 0.12 (0.01) 0.05 (0.01)
Table 9: Power of the Levenshtein MMD test to distinguish the fp32 model from alternatives in $L=50$ token space at $N=10m$ samples, where $m=25$ is the number of prompts in the prompt distribution. Averages and standard errors are reported over ten prompt distributions. The column represents the alternative distribution $Q$. The FPR (fp32) column represents the false positive rate when comparing the full-precision model to itself.
C.2 Additional results from §4.3
[…]
Discussion: power decreases in higher-dimensional space.
Table 10 shows the power of the Hamming MMD test to distinguish between pairs of distributions $(P,Q)$ in $L=$1000$$ character space. As expected, moving to this higher-dimensional space decreases the power of the test; the powers in the last three rows are lower than those in Table 8 for the same models in $L=50$ token space. These power losses can be compensated for by increasing the sample size: Table 11 repeats the experiments in Table 10 with $N=50m$ instead of $N=10m$. If we were to use this sample size in §5, each audit would still cost $<$ \$5.
Note that even in the reduced power setting, model swaps are easy to detect (Table 10, top rows).
model Mistral 7B Llama-3 8B Llama-3.1 8B Llama-3 70B Llama-3.1 70B
alternative
Mistral 7B 0.06 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Llama-3 8B 1.00 (0.00) 0.05 (0.01) 0.76 (0.00) 0.98 (0.00) 0.95 (0.00)
Llama-3.1 8B 1.00 (0.00) 0.83 (0.05) 0.07 (0.01) 1.00 (0.00) 0.53 (0.00)
Llama-3 70B 1.00 (0.00) 0.99 (0.00) 1.00 (0.00) 0.07 (0.01) 0.89 (0.00)
Llama-3.1 70B 1.00 (0.00) 0.98 (0.00) 0.75 (0.07) 0.99 (0.00) 0.06 (0.01)
Phi-3 mini 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Gemma-2 9B 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
OLMo 7B 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
gpt-4o-mini 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
gpt-4o 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
gpt-3.5-turbo-0125 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
gpt-3.5-turbo-1106 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
gpt-4-0125-preview 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
watermark 0.23 (0.02) 0.62 (0.05) 0.32 (0.04) 0.57 (0.06) 0.26 (0.03)
int8 0.15 (0.02) 0.30 (0.06) 0.07 (0.01) 1.00 (0.00) 0.99 (0.01)
nf4 0.44 (0.05) 0.38 (0.05) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
@@
Table 10: Power of the Hamming MMD test to distinguish between pairs of models (and other alternatives) in $L=$1000$$ character space at $N=10m$, where $m=25$ is the number of prompts covered by $\pi$. The column represents the null distribution $P$. Italicized entries are FPRs. Note that unlike §4.1, the test is now run in higher-dimensional character space, which generally reduces power.
Sample size watermark nf4 int8 FPR (fp32)
$N=10m$ 0.62 (0.05) 0.38 (0.05) 0.30 (0.06) 0.05 (0.01)
$N=50m$ 1.00 (0.00) 1.00 (0.01) 0.96 (0.02) 0.07 (0.01)
Table 11: Power of the Hamming MMD test to distinguish between the fp32 Llama-3 8B and quantized or watermarked alternatives in $L=$1000$$ character space. We compare power at $N=10m$ and $N=50m$, where $m=25$ is the number of prompts covered by $\pi$. Increasing the sample size can compensate for losses of power.
C.3 Detecting incorrect decoding parameters
Our problem formulation in §2 compares two distributions in general: this includes cases where the API $Q$ samples from the same model as the reference distribution $P$, but with different decoding parameters (e.g., temperature $\tau$). Table 12 shows that detecting differences in the temperature parameter is quite easy; this parameter significantly affects the distribution of completions.
Model $\tau=0.5$ $\tau=1.5$
Llama-3 8B 1.00 (0.00) 1.00 (0.00)
Llama-3.1 8B 1.00 (0.00) 1.00 (0.00)
Mistral 8B 1.00 (0.00) 0.97 (0.02)
Table 12: Power of the Hamming MMD test to distinguish between models sampled at temperature $\tau=1$ ($P$) vs. other temperatures ($Q$). This decoding parameter mismatch can be detected with near-perfect power.
C.4 Effect of the prompt distribution
In §4, we evaluated tests on a prompt distribution $\pi$ drawn from a Wikipedia language modeling task and supported on $m=25$ prompts. In this appendix, we extend these evaluations to different prompt distributions.
Effect of the number of prompts $m$.
@@ C.4 Effect of the prompt distribution
In Figure 8, we continue to use the Wikipedia task, but vary the number of prompts $m$ in the prompt distribution. Larger $m$ values increase the power of the test, suggesting that users benefit from testing many prompts together, so long as the sample size is increased proportionally.
@@
Effect of the task.
We explored whether the “open-endedness” of the task affects power: intuitively, one might expect that more creative tasks lead to completion distributions that are higher entropy, which might make testing harder. To test this, we evaluated power across several prompt distributions: on the most constrained side of the spectrum, we experimented with HumanEval (code, Box B.1); on the most creative side, UltraChat (chatbot dialogues, Box B.1), and in the middle, language modeling with Wikipedia (Box 4). We additionally experimented with tightly concentrated prompts (just Wikipedia in English, just Wikipedia in French, etc.) vs. diverse prompt distributions (Wikipedia mixed with HumanEval and UltraChat).
Table 13 lists the Hamming MMD’s power against the local alternatives (nf4, int8, watermark) at $N=10m$ with tests conducted in token space ($L=50$ token completions). We make several observations:
  • We do not observe that creative tasks are harder to test than constrained tasks: if this were the case, we would expect the power to be lowest for UltraChat and highest for HumanEval across all alternatives.
  • We also do not observe that diverse prompt distributions are harder to test than concentrated prompt distributions: if this were the case, we would expect the power to be higher for the single-language Wikipedia prompt distributions than the mixed-task prompt distributions.
  • The prompt distribution mainly affects the difficulty of detecting quantization, not watermarking.
  • For comparison, we also provide the results of the one-hot kernel in Table 14. The Hamming kernel consistently outperforms the one-hot baseline on all prompt distributions, matching our main conclusion of §4.1 about the relative strength of kernels.
$m$ watermark nf4 int8 FPR (fp32)
HumanEval (constrained) 20 0.99 (0.00) 1.00 (0.00) 0.32 (0.00) 0.06 (0.00)
Wikipedia (language modeling) English only 25 0.98 (0.01) 0.49 (0.04) 0.11 (0.01) 0.06 (0.01)
French only 25 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 0.06 (0.01)
German only 25 0.98 (0.01) 0.71 (0.08) 0.27 (0.06) 0.05 (0.00)
Spanish only 25 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 0.05 (0.01)
Russian only 25 0.98 (0.01) 0.71 (0.05) 0.24 (0.04) 0.05 (0.01)
All languages 25 1.00 (0.00) 0.93 (0.04) 0.79 (0.08) 0.05 (0.01)
UltraChat (creative) 20 0.89 (0.00) 1.00 (0.00) 0.04 (0.00) 0.04 (0.00)
Wikipedia + UltraChat 25 0.81 (0.05) 1.00 (0.00) 0.10 (0.02) 0.06 (0.01)
Wikipedia + UltraChat + HumanEval 25 0.95 (0.01) 1.00 (0.00) 0.35 (0.05) 0.04 (0.01)
Table 13: Hamming MMD’s power against the local alternatives (nf4, int8, watermark) at $N=10m$ with tests conducted in token space ($L=50$ token completions).
$m$ watermark nf4 int8 FPR (fp32)
HumanEval (constrained) 20 0.46 (0.00) 0.77 (0.00) 0.16 (0.00) 0.05 (0.00)
Wikipedia (language modeling) English only 25 0.06 (0.01) 0.05 (0.01) 0.04 (0.01) 0.01 (0.00)
French only 25 0.04 (0.01) 0.06 (0.01) 0.05 (0.00) 0.01 (0.00)
German only 25 0.04 (0.01) 0.04 (0.01) 0.04 (0.01) 0.01 (0.00)
Spanish only 25 0.04 (0.01) 0.06 (0.01) 0.04 (0.01) 0.02 (0.00)
Russian only 25 0.04 (0.01) 0.05 (0.01) 0.03 (0.01) 0.01 (0.00)
All languages 25 0.04 (0.00) 0.05 (0.01) 0.04 (0.01) 0.01 (0.00)
UltraChat (creative) 20 0.07 (0.00) 0.48 (0.00) 0.04 (0.00) 0.01 (0.00)
Wikipedia + UltraChat 25 0.09 (0.01) 0.05 (0.01) 0.07 (0.01) 0.04 (0.01)
Wikipedia + UltraChat + HumanEval 25 0.22 (0.02) 0.32 (0.03) 0.05 (0.01) 0.05 (0.01)
Table 14: One-hot MMD’s power against the local alternatives (nf4, int8, watermark) at $N=10m$ with tests conducted in token space ($L=50$ token completions).
C.5 comparing two-sample and approximate goodness-of-fit tests
@@
C.6 asymmetric sampling costs